### v2

# THE MORE, THE MERRIER: CONTRASTIVE FUSION FOR HIGHER-ORDER MULTIMODAL ALIGNMENT

Abstract
Learning joint representations across multiple modalities remains a central challenge in multimodal machine learning. Prevailing approaches predominantly operate in pairwise settings, aligning two modalities at a time. While some recent methods aim to capture higher-order interactions among multiple modalities, they often overlook or insufficiently preserve pairwise relationships, limiting their effectiveness on single-modality tasks. In this work, we introduce Contrastive Fusion (ConFu), a framework that jointly embeds both individual modalities and their fused combinations into a unified representation space, where modalities and their fused counterparts are aligned. ConFu extends traditional pairwise contrastive objectives with an additional fused-modality contrastive term, encouraging the joint embedding of modality pairs with a third modality. This formulation enables ConFu to capture higher-order dependencies, such as XOR-like relationships, that cannot be recovered through pairwise alignment alone, while still maintaining strong pairwise correspondence. We evaluate ConFu on synthetic and real-world multimodal benchmarks, assessing its ability to exploit cross-modal complementarity, capture higher-order dependencies, and scale with increasing multimodal complexity. Across these settings, ConFu demonstrates competitive performance on retrieval and classification tasks, while supporting unified one-to-one and two-to-one retrieval within a single contrastive framework. We release our code and dataset at https://github.com/estafons/confu.
1 Introduction
Learning joint representations across multiple modalities is a fundamental goal in multimodal machine learning [44]. Contrastive frameworks such as CLIP [26] and ALIGN [16] have demonstrated that aligning paired modalities, such as images and text, enables powerful zero-shot transfer and retrieval. However, these approaches remain intrinsically pairwise, capturing only correlations between two modalities while overlooking the higher-order dependencies that emerge when three or more modalities interact [11, 16, 10, 48, 35, 34, 47, 41, 42, 36, 46]. Beyond this limitation, multimodal models often exhibit modality competition [14], where one modality dominates and suppresses others when signal strength or reliability differs. For instance, strong visual cues may overshadow weaker auditory ones, biasing the joint representation toward a single source. Yet many real-world settings require reasoning over complementary cues: a song emerges from melody and lyrics, a 3D design from sketch and text, and a scene from both visual and auditory signals. These challenges motivate a broader question:
To address this, we propose Contrastive Fusion (ConFu), a framework that unifies pairwise and higher-order multimodal supervision within a single contrastive objective. ConFu constructs fused representations of all subsets of modalities and aligns all combinations of unimodal and fused representations (Fig. 2). Unlike previous methods that primarily target specific one-to-one (11) [11, 10], two-to-one (21) [27, 5], or both within predominantly text-centered objectives [6, 4], ConFu accounts for all pairwise and higher-order modality combinations within a unified objective (Fig. 1).
Evaluating such models is further complicated by the scarcity of publicly available datasets that naturally include three or more co-occurring modalities. To address this, we construct a bird-centric dataset where image, audio, and text provide complementary but related information. This setting enables a systematic examination of higher-order alignment under zero-shot and few-shot regimes. Overall, our main contributions are as follows:
We introduce a unified contrastive framework that generalizes CLIP-like pairwise multimodal alignment to capture higher-order dependencies among three modalities.
We construct Bird-MML, a synthetic dataset of artificial triplets that serves as a pretraining resource to evaluate whether models capture multimodal complementarity.
We demonstrate that ConFu achieves competitive performance in retrieval and classification tasks, supporting both one-to-one and two-to-one retrieval within a single contrastive framework.
2 Related Work
Jointly-aligned models. Methods such as AudioCLIP [11], VALOR [3], and VAST [4] extend CLIP-style contrastive learning to three or more modalities by training on datasets where all signals co-occur. AudioCLIP employs independent contrastive objectives for each modality pair, while VALOR introduces a multimodal grouping alignment loss coupled with caption generation to connect multiple modalities through text. VAST combines contrastive and generative losses to unify video, audio, and subtitles in a shared representation. Other architectures such as mPLUG-2 [40], UMT [22], and the Everything-at-Once Fusion Transformer [21] employ modular decoders or fusion transformers for multimodal understanding and generation. While these methods learn representations across multiple modalities, their objectives are typically limited to pairwise alignments or remain primarily text-centric, with broader multimodal relationships emerging implicitly from shared training signals rather than through explicit higher-order alignment.
Pivot-aligned models. An alternative strategy uses a pivot modality as a common reference space. Each non-pivot modality is aligned to this anchor, enabling scalable training without requiring full multimodal co-occurrence. For example, ImageBind [10] aligns diverse modalities (audio, depth, IMU) to an image-based embedding space, while LanguageBind [48] uses text as the pivot. Later variants such as OmniBind [35], FreeBind [34], and Ex-MCR [46] extend this paradigm through pseudo-pair generation, offering an alternate data-efficient mechanism. Similarly, Uni3D [47], ULIP [41], and ULIP-2 [42] project new modalities into a frozen CLIP embedding space,mapping new modalities to CLIP’s pretrained space without updating its parameters. While this pivot-based design improves scalability, it remains inherently pairwise: all interactions are mediated through the pivot, preventing direct modeling of dependencies among non-pivot modalities. As a result, the learned embedding space captures independent pairwise relations rather than true higher-order structure.
Modeling Higher-Order Dependencies. Recent work in self-supervised representation learning has attempted to model relationships between more than two modalities through specialized loss formulations. Symile [27] leverages total correlation to integrate complementary cues, TRIANGLE [5] uses the area spanned by three embedding vectors as a measure of similarity, and GRAM [6] measures cross-modal similarity for modalities through the Gramian volume spanned by their vectors. While these methods jointly model multiple modalities, Symile and TRIANGLE require all modalities to be present at inference time, making them incompatible with standard 11 retrieval.
In contrast, our framework explicitly models both pairwise and higher-order alignments within a single contrastive objective. It provides a simple, architecture-agnostic formulation that unifies pairwise and higher-order multimodal interactions, enabling scalable and expressive cross-modal representation learning, with the only additional computational cost coming from lightweight MLP layers.
3 Method
3.1 Proposed Multimodal Contrastive Framework
We propose a new framework for contrastive learning with modalities (Fig. 2) that unifies pairwise and higher-order contrastive objectives into a single multimodal training process that maximizes a lower bound on the total correlation among modalities (Sec. 3.2).
Each modality is mapped into a feature space by a modality-specific encoder and then projected into a shared latent space by a projector :
The encoder parameters and projector parameters are learned jointly through contrastive optimization.
Pairwise Contrastive Objectives ().
For each modality pair , we estimate their MI via the InfoNCE lower bound Eq. 2. Let denote a density ratio function parameterized by , which models the likelihood of joint co-occurrence over independence. Given a minibatch of paired samples , the empirical loss is
Summing over all ordered modality pairs yields
Higher-Order Contrastive Objectives ().
To model higher-order dependencies , we introduce a fusion network with encoded features:
where are trainable fusion parameters. In our experiments, we have maintained to be a small-scale network (e.g., a shallow MLP) to capture cross-modal interactions without significantly increasing model complexity. Each fused representation is then aligned with the remaining modality , using the ratio density estimation function parameterized by :
In practice, both density ratio estimators in Eq. 2 and in Eq. 5 are implemented as temp.-scaled dot-product similarities. The higher-order objective aggregates all such triplets:
Combined Objective.
The total training loss jointly maximizes both pairwise and fused MI lower bounds:
where controls the relative contribution of pairwise versus higher-order supervision (See Appendix B.1 for analysis of ’s impact). Minimizing Eq. 7 corresponds to maximizing the contrastive lower bound on (Eq. 13), thus promoting both (i) pairwise alignment and (ii) cross-modal consistency through fused interactions.
3.2 Theoretical Motivation
The total correlation (TC) [37], also known as multi-information [28], quantifies the joint statistical dependency among multiple random variables as the Kullback–Leibler divergence between their joint distribution and the product of marginals, . Equivalently,
and vanishes if and only if the variables are mutually independent (). Using the chain rule of MI, it decomposes as
Averaging over all permutations yields the symmetric form
where denotes the set of all six permutations of the index triplet . This interpretation of TC separates pairwise and higher-order dependencies.
Contrastive Lower Bound.
Each MI term in Eq. 10 is approximated using the InfoNCE lower bounds [23]. Specifically, the pairwise and higher-order terms satisfy
Substituting these inequalities into Eq. 10 yields the tractable contrastive lower bound on ,
so that maximizing total correlation reduces to minimizing the InfoNCE losses across all pairwise and higher-order relations. When each density ratio function or is sufficiently expressive and approximates the optimal InfoNCE critic [23] (i.e., the log–density ratio between the joint and the product of marginals) the bound becomes tight as , recovering the true .
Interpretation.
3.3 Compare with other higher-order methods
Recent approaches such as Symile, GRAM, and TRIANGLE adopt a single joint alignment objective across all modalities, without separating pairwise from higher-order dependencies. Symile maximizes a total-correlation bound using a multilinear similarity , GRAM enforces global alignment via the parallelotope volume induced by the Gram matrix, and TRIANGLE specializes this idea by introducing a triangle-area–based measure of three-way similarity. In contrast, our method decomposes total correlation into interpretable information-theoretic factors, averaging over modality permutations the sum of pairwise and higher-order dependencies, thus preserving holistic expressivity while enabling structured optimization and reliable performance with any subset of modalities. Unlike prior approaches that use a single critic to implicitly model dependencies, we factorize them at the loss level, enabling independent control of pairwise and higher-order supervision while preserving InfoNCE stability and achieving balanced performance within a single model.
3.4 Motivating higher-order dependencies
To illustrate the reasons that modeling higher-order (synergistic) dependencies is crucial, we consider the XOR task introduced in Symile. Three binary variables , , and are sampled with , , and the relationship:
The task is to predict the representation from the pair . The parameter controls the level of synergy: for , is independent of , while increasing raises the conditional MI and without introducing any pairwise dependence (). Thus, only models capturing joint interactions, rather than pairwise correlations, can solve the task. Zero-shot retrieval at varying levels of (Fig. 3) confirms that ConFu and Symile can successfully solve the task.
4 Bird-MML: A Multimodal Dataset for Audio-Visual Complementarity
Despite substantial progress in multimodal learning, there remains a lack of standardized and publicly available realistic datasets that capture cross-modal relationships that are suitable for evaluating both pairwise and higher-order multimodal alignment. Bird species provide an ideal testbed for such multimodal evaluation, as their identity and behavior are naturally expressed through both visual and acoustic cues. In many cases, fine-grained recognition further depends on complementary information such as visual appearance for morphology and audio for vocalization, making the domain well suited for studying multimodal interaction.
Existing Bird Datasets. SSW60 [30] offers unaligned images and audio as well as video for 60 species, VB100 [9] focuses on fine-grained video recognition for 100 species. Moreover, iNatSounds [2] contains large-scale bioacoustic recordings without visual counterparts, and lastly, BioTrove-Train [43] includes millions of images across diverse species with uneven per-class distributions. Although those datasets provide valuable resources, a joint combination of images, audio, and text in a form suitable for multimodal pretraining and evaluation is missing.
Dataset Overview. To bridge this gap, we construct Bird-MML, a synthetic multimodal bird dataset integrating artificially paired images, audio, and text for 150 species, including all those present in SSW60 and VB100. The dataset is designed for scalable pretraining and cross-modal learning, with balanced representation across species. A detailed overview of the data generation pipeline is provided in Appendix D.1. Images are sourced from the iNaturalist open dataset, restricted to research-grade, Creative Commons–licensed observations. For each species, 1,000 samples are randomly selected.
Audio and Text Generation. Audio recordings are collected from Xeno-Canto [39], segmented into 10-second clips, and zero-padded when necessary. In cases of limited availability, existing clips are reused to maintain uniform sample counts. Metadata such as call type, sex, and life stage are preserved. Text descriptions are generated by combining three complementary sources: image captions from InstructBLIP2 [7], audio metadata, and short summaries from corresponding Wikipedia entries. These elements are fused using google/gemma-2-2b-it [29] to yield a single multimodal caption per instance.
Dataset Characteristics. While generated captions may contain minor factual noise, they effectively capture species-relevant attributes and contextual cues useful for representation learning. The resulting dataset contains 149,681 triplets (image, audio, text), each grounded by species identity and visual–acoustic correspondence. Approximately 43% of species required limited audio reuse due to data scarcity. Examples of artificially aligned triplets are shown in Appendix D.1.The dataset maintains balanced per-species coverage and is well-suited for multimodal pretraining. Downstream evaluations are performed on naturally paired video-audio data to assess real-world cross-modal performance.
Ethical Considerations. All data were collected from publicly available sources under Creative Commons licenses. Ownership and attribution remain with original contributors. Automatically generated captions may contain minor factual errors or biases and should not be interpreted as verified biological information. The dataset is released solely for non-commercial research use.
| Method | A | V | A+V |
| CLIP | 41.1 0.3 | 63.0 0.4 | - |
| Tri-CLIP | 40.0 0.4 | 62.5 0.4 | - |
| Symile [27] | - | - | 70.9 0.3 |
| GRAM [6] | 9.8 0.4 | 63.9 0.8 | 64.4 0.5 |
| TRIANGLE [5] | - | - | 64.9 0.3 |
| ConFu | 39.7 0.1 | 64.6 0.3 | 71.2 0.2 |
5 Experimental Results
Our evaluation covers multimodal benchmarks across biodiversity (SSW60 [30], VB100 [9]), affective (MOSI [45], UR-FUNNY [12], and MUStARD [1] included in MultiBench [20]), and multimedia (AV-MNIST [31]) domains, encompassing both real-world and synthetic datasets.
These datasets present different levels of cross-modal difficulty, enabling holistic evaluation of alignment, representation quality, and scalability. We compare against representative contrastive and higher-order multimodal approaches: (i) a Bi-modal CLIP (referenced as CLIP from now on) trained with standard two-modality contrastive loss, (ii) a Tri-modal CLIP (referenced as Tri-CLIP from now on) trained with pairwise objectives across all modalities, (iii) Symile [27] which models higher-order dependencies via total correlation, (iv) TRIANGLE [5] and GRAM [6], which use geometric similarity measures among 2 or more modalities. Experiments cover (i) zero-shot classification, (ii) 11 and 21 retrieval, and (iii) linear-probe classification under few-shot and full-data regimes. Details about the experimental setup and evaluation can be found in Appendix A.
5.1 AV-MNIST
We evaluate our framework on the AV-MNIST dataset, which contains paired audio spectrograms augmented with randomly sampled background noise from ESC-50 [25] and images of MNIST digits with degraded visual features. We additionally introduce a text modality derived from class labels using natural language templates such as ‘the digit is a {class}”, ‘a photo of a {class}”, and ‘a picture of a {class}”. As provided in Table 1, our approach achieves the highest performance across all baselines, with Symile trailing closely behind. Fusing audio and visual inputs yields an 8% improvement over the strongest unimodal baseline, demonstrating the benefit of complementary multimodal supervision. Moreover, even when evaluated on a single modality (vision), our model surpasses the corresponding unimodal baseline by 1.5%, indicating that multimodal training enhances individual modality representations.
| Target | M1 | M2 | M3 | Mean | ||||||||
| Query | M2 | M3 | M23 | M1 | M3 | M13 | M1 | M2 | M12 | 1→1 | 2→1 | |
| MOSI | Trimodal CLIP | 22.9 1.9 | 20.5 1.6 | – | 23.5 3.6 | 23.3 1.8 | – | 20.9 2.0 | 24.3 2.1 | – | 22.6 | – |
| Symile [27] | - | - | 16.3 0.9 | - | - | 18.1 2.4 | - | - | 17.1 1.7 | – | 17.2 | |
| GRAM [6] | 6.0 3.1 | 10.3 2.9 | 16.3 2.5 | 7.7 2.3 | 0.3 0.6 | 7.9 3.4 | 12.2 4.1 | 0.2 0.2 | 12.0 4.2 | 6.1 | 12.1 | |
| TRIANGLE [5] | - | - | 8.3 1.3 | - | - | 4.9 0.7 | - | - | 8.8 1.4 | – | 7.3 | |
[…]
| UR-FUNNY | Trimodal CLIP | 3.7 0.6 | 4.0 0.6 | – | 3.8 0.5 | 16.2 1.6 | – | 4.0 0.4 | 16.8 1.2 | – | 8.1 | – |
| Symile [27] | - | - | 3.7 0.5 | - | - | 15.4 0.9 | - | - | 16.5 0.6 | – | 11.9 | |
| GRAM [6] | 3.2 0.5 | 3.0 0.9 | 3.9 0.8 | 3.2 0.9 | 0.1 0.2 | 3.1 0.4 | 3.9 0.4 | 0.1 0.0 | 3.3 0.5 | 2.3 | 3.4 | |
| TRIANGLE [5] | - | - | 4.2 0.4 | - | - | 3.3 0.7 | - | - | 3.6 0.9 | – | 3.7 | |
[…]
| MUStARD | Trimodal CLIP | 70.7 4.1 | 29.1 3.6 | – | 70.5 4.3 | 26.5 3.2 | – | 30.4 3.4 | 24.6 2.8 | – | 42.0 | – |
| Symile [27] | - | - | 61.5 6.3 | - | - | 57.0 3.0 | - | - | 21.3 4.4 | – | 46.6 | |
| GRAM [6] | 61.7 8.0 | 24.8 5.3 | 81.0 4.1 | 67.2 5.4 | 24.8 5.3 | 67.8 5.4 | 27.9 8.7 | 5.1 2.1 | 31.2 9.2 | 35.2 | 60.0 | |
| TRIANGLE [5] | - | - | 68.7 4.2 | - | - | 59.9 3.7 | - | - | 21.9 5.0 | – | 50.2 | |
[…]
5.2 Affective Computing Benchmarks
We next evaluate on the MOSI, UR-FUNNY, and MUStARD datasets, which involve text, audio, and video for multimodal affect understanding. We assess (i) 11 retrieval, (ii) 21 retrieval, and (iii) multimodal classification via a linear probe. Table 2 shows that ConFu consistently improves cross-modal retrieval performance. In particular, the 21 retrieval setting surpasses 11 retrieval on UR-FUNNY and MUStARD, suggesting that combining two input modalities provides richer and more discriminative cues.
Notably, GRAM, the only baseline supporting both retrieval configurations, struggles with 11 alignments in general and fails for modalities M2 and M3 on MOSI and UR-FUNNY, suggesting that balancing 21 and 11 alignment is non-trivial. Furthermore, methods such as TRIANGLE and Symile, which focus on trimodal alignment without explicitly modeling pairwise relations, perform consistently worse than ConFu. Overall, ConFu ranks first or second across nearly all querytarget retrieval settings, in both 11 and 21 configurations, demonstrating robust and consistent cross-modal alignment. Moreover, Table 3 reports linear-probe classification results, showing that ConFu attains the highest accuracy on UR-FUNNY and remains competitive on MUStARD and MOSI.
5.3 Fine-Grained Bird Classification
ConFu is further evaluated on fine-grained bird classification using the VB100 and SSW60 datasets. Models are pretrained on Bird-MML, our curated dataset of 150K audio–image–text triplets and assessed under zero-shot and few-shot regimes. We consider two evaluation settings: (1) the multi-frame protocol from [30], where each video clip is divided into eight segments, the middle frame of each segment is sampled, and the mean embedding of the eight frames is used as the visual representation; and (2) a single-frame setting, providing a more constrained visual input per video. Table 4 reports zero-shot results in the multi-frame setting. ConFu achieves the highest accuracy on SSW60 (71.4%), confirming that multimodal fusion improves fine-grained discrimination when both modalities are informative. On VB100, where the audio modality is largely uninformative ( unimodal accuracy), the fused model (A+V) maintains performance close to the best visual-only baseline (18.15% vs. 20.69%), demonstrating robustness to distracting modalities. In contrast, GRAM, Symile, and TRIANGLE exhibit a notable decline in accuracy (5–10%) relative to ConFu. Few-shot trends in Fig. 4 mirror these findings. Fusion offers clear advantages on SSW60 and stable performance on VB100, showing that ConFu generalizes well even when one modality is weak.
| SSW60 Acc. (%) | VB100 Acc. (%) | |||||
| Method | A | V | A+V | A | V | A+V |
| CLIP | 29.9 | 60.3 3.4 | – | 4.2 | 18.2 0.5 | – |
| Tri-CLIP | 31.1 | 59.7 3.2 | – | 3.9 | 17.7 0.5 | – |
| Symile [27] | - | - | 61.4 0.2 | - | - | 13.2 0.4 |
| TRIANGLE [5] | - | - | 58.6 3.3 | - | - | 8.7 0.4 |
| GRAM [6] | 0.7 | 59.0 3.3 | 54.9 3.2 | 1.3 | 12.9 0.4 | 7.6 0.3 |
| ConFu | 30.3 | 59.8 3.3 | 65.5 2.5 | 3.4 | 16.7 0.5 | 16.4 0.4 |
In the single-frame setting, designed to evaluate performance with limited visual information, we averaged classification results across 19 uniformly sampled frames to reduce bias. As provided in Table 5, the fused (A+V) model exceeds the best baseline by roughly 4% on SSW60 and complementary audio–visual cues when visual input is sparse. On VB100, performance drops slightly (by 2–3%) due to the weaker audio modality ( accuracy). Overall, ConFu benefits from richer multimodal supervision, maintaining strong robustness under modality degradation while preserving scalability and simplicity. Additional results on a few-shot adaptation are in the Appendix C.1.
6 Ablation
Modality competition.
Although multimodal learning seeks to combine complementary cues, models often show modality dominance, where stronger signals suppress weaker ones [14]. To assess how our fusion is affected by this effect, we analyze in Fig. 5 the overlap of prediction errors across modalities and fusion. Results reveal visual dominance: The larger portion of correct samples (39%) are predicted by both the fusion and vision models, while the datapoints that only audio predicts correctly or fusion-audio overlap remain small (5–6%), and purely audio-driven successes are rare (2%). Nevertheless, about 5% of samples are uniquely solved by the audiovisual model, indicating that fusion captures complementary information beyond either modality alone. Finally, some classes benefit substantially from fusion (e.g., Label 31 see Appendix B.5), while others remain uniformly difficult.
Ablation on Noise-Induced Distribution Shift.
To examine the sensitivity of each model to distribution shift, we introduced Gaussian noise to the image modalitiy only during testing, with a standard deviation of 0.05 corresponding to roughly 20dB SNR. This perturbation simulates a mismatch between the training and test distributions without retraining the models. As shown in Table 6, ConFu achieves the highest accuracy (45.4%) under these conditions. In contrast, the unimodal baseline (Tri-CLIP), exhibits a sharper performance drop, while Symile shows moderate robustness. Interestingly, TRIANGLE and GRAM perform even worse compared to Tri-CLIP under visual degradation. Overall, these results indicate that ConFu’s fusion strategy improves stability under distribution shifts affecting the image modality. Regarding more experiments in noise induced distribution shifts see Appendix B.4.
7 Discussion
Empirical results across synthetic and real-world datasets demonstrate that ConFu effectively captures complementary cross-modal structure. On the synthetic XOR benchmark, ConFu successfully models the underlying multimodal dependencies that pairwise and geometric baselines miss even with higher-dimensional representations (Appendix B.2). In AV-MNIST, ConFu dominates unimodal approaches in zero-shot classification, indicating successful utilization of both modalities. The 21 retrieval evaluations on UR-FUNNY, and MUStARD further confirm that combining two query modalities provides richer semantic grounding than either considered in isolation. An interesting observation in the 21 case on the MOSI dataset is that fusion slightly underperforms compared to the pairwise baseline. Remarkably, this issue is substantially alleviated by introducing partial modality masking (Appendix B.3), supporting the hypothesis that suboptimal performance stems from shortcut learning in cases where redundant information dominates.
In fine-grained bird classification, ConFu maintains or slightly exceeds unimodal baselines. On SSW60, where both modalities contribute meaningfully, fusion achieves the best accuracy, on VB100, where acoustic information is limited, performance remains close to the visual baseline. Moreover, modality overlap analyses indicates that a subset of samples is correctly resolved only when both modalities are utilized, reflecting the model’s ability to integrate complementary evidence. Finally, under visual perturbations, ConFu degrades more gracefully, outperforming all baselines by a clear margin. SYMILE ranks second, remaining above unimodal performance, while TRIANGLE and GRAM drop below the single-modality baseline.
Limitations and Future Work. ConFu is well-positioned to scale to any number of modalities, offering flexibility for diverse multimodal tasks. However, as the number of modalities increases, computational demands may also rise. Depending on the task, selectively pruning certain alignment terms, shown in Appendix E, can help maintain efficiency. For example, omitting many-to-many alignments, when assuming that only single-modality queries or targets will be needed, opens a promising avenue for future adaptation and research.
ConFu currently relies on fully aligned modalities during training, and exploring ways to relax or bypass this requirement remains crucial for many applications. In particular, enabling ConFu to handle partially missing modalities, such as training with incomplete triplets or drawing on ideas related to pseudopair construction [35, 46, 36], could potentially make the framework more robust and broadly applicable across real-world multimodal settings. Further prospective directions could include extending ConFu to incorporate factorized representations inspired by recent factorization-based and redundancy-reduction approaches [19, 33, 8], developing adaptive mechanisms to regulate modality competition [18, 17, 24, 13, 38], and investigating how ConFu can enhance robustness in multimodal retrieval, particularly when one or more modalities are noisy, incomplete, or severely corrupted.
8 Conclusion
We presented ConFu, a unified contrastive framework that jointly optimizes pairwise and higher-order alignments to capture pairwise and synergistic relationships across modalities. The method supports both 1→1 and 2→1 retrieval, improving multimodal alignment and robustness without modifying encoder architectures. Overall, ConFu demonstrates strong adaptability across diverse datasets, domains, and challenging conditions such as noise-induced distribution shifts, modality degradation, and missing modalities, while introducing only minimal overhead through lightweight MLPs atop the encoders, establishing it as a robust and scalable solution for multimodal alignment in real-world applications. The accompanying Bird-MML dataset establishes a controlled setup for studying multimodal complementarity under realistic audio–visual conditions.
9 Acknowledgments
This work was supported by the HARSH project (project no. Y3TA - 0560901) within the framework of the National Recovery and Resilience Plan–Greece 2.0–with funding from the European Union–NextGenerationEU; the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I.’s Research Projects to Support Faculty Members & Researchers” (H.F.R.I. Project Number: 26302); by the TITAN ERA Chair project (contract no. 101086741) within the Horizon Europe Framework Program of the European Commission; the METHUSALEM program (METH/26/003, “Methusalem-BioMedAI”); the Flemish Government AI Research Program (Leuven.AI – KU Leuven); the FWO project “Task- and Device-Agnostic Artificial Intelligence for EEG Analysis” (G046925N); and the Horizon Europe project “AI-PROGNOSIS” (Grant Agreement No. 101080581).
Supplementary Material
Supplementary Material
This supplementary document provides additional details and results that complement the main paper. It is organized as follows:
Experimental Details. We present comprehensive descriptions of model configurations, hyperparameters, training procedures, and evaluation protocols to ensure reproducibility.
Additional Ablation Studies. We include supplementary ablation experiments that highlight additional aspects of this work, examining how different model components and design choices influence overall performance.
Additional Results. This section includes further quantitative results that extend and support the findings presented in the main paper.
Datasets. Additional information regarding the datasets used and our new dataset Bird-MML.
Appendix A Experimental Details
A.1 Experiments on Bird-MML
Dataset preprocessing
For the audio modality, Each audio clip was truncated or zero-padded to a duration of 10 seconds and resampled to 22.05 kHz. Mel spectrograms were computed using a window size of 1024 and a hop length of 512, resulting in 128 Mel frequency bands. Each spectrogram was treated as a single-channel image to ensure compatibility with a ResNet-50 backbone.
For the image modality, Images were resized to pixels and normalized. No data augmentation was applied.
Model Architecture.
The multimodal architecture employed is composed of three modality-specific encoders:
Text encoder: BERT transformer.
Image encoder: ResNet-50.
Audio encoder: ResNet-50 applied to Mel-spectrogram inputs.
Each encoder outputs a 512-dimensional embedding. A two-layer multilayer perceptron (MLP) was used for multimodal feature fusion. All models and baselines were trained using identical hyperparameters, without any additional hyperparameter search or tuning.
Training used the AdamW optimizer with a learning rate of and a weight decay of . A batch size of 64 was used throughout the training process. The learning rate followed a cosine annealing schedule with a minimum of . The fusion loss coefficient was set to .
A.2 MOSI, MUStARD, and UR-FUNNY
For these datasets, all models were trained for 100 epochs, producing 256-dimensional representations. We used a batch size of 128 and a learning rate of . Following prior work [19], each modality encoder was implemented as a 5-layer Transformer with 5 attention heads.
Fusion was achieved through a 2-layer MLP that combined unimodal embeddings into a single fused representation. Dataset splits followed the official MultiBench configuration for training, validation, and testing.
A.3 AVMNIST
For the text modality, we used a BERT transformer. Image, spectrogram, and satellite-view inputs were encoded using ResNet-18 backbones initialized from ImageNet and fully trained. Spectrograms provided by the MultiBench dataloaders were used without modification.
Embeddings had a 256-dimensional size. Dataset splits again followed the MultiBench protocol.
Training Setup.
All models were trained from scratch for 30 epochs using the AdamW optimizer with a learning rate of and a batch size of 256. The best model was selected based on validation loss. Each experiment was repeated with five random seeds, and we report the mean and standard deviation across runs. For our model, MLP-based fusion was used with a balancing parameter between pairwise and fusion losses.
A.4 XOR Experiment
For the XOR experiment, we mapped all modalities to embeddings of size 128. The dataset consisted of 10,000 training samples and 5,000 test samples. Models were trained for 50 epochs with a batch size of 512 and a learning rate of .
Each encoder was implemented as a 2-layer MLP, and we used a balanced fusion coefficient of .
Appendix B Additional Ablation Studies
B.1 Effect of the Parameter
As shown in Table 7, performance is not overly sensitive to . Accuracy remains stable while recall varies smoothly, indicating a controlled trade-off with minimal tuning required. Identifying a principled way to set or eliminate is an important direction.
| Dataset | Acc. (%) | |||
| MOSI | 0.2 | 67.7 1.7 | 19.5 1.0 | 20.1 1.9 |
| 0.5 | 66.7 2.1 | 19.5 1.5 | 19.6 2.0 | |
| 0.8 | 65.8 2.0 | 13.5 2.2 | 15.6 1.6 | |
| UR-Funny | 0.2 | 64.2 0.7 | 8.0 0.4 | 14.0 0.9 |
| 0.5 | 64.9 1.0 | 7.4 0.2 | 13.6 0.6 | |
| 0.8 | 64.6 0.7 | 6.9 0.3 | 13.2 0.7 | |
| MUSTARD | 0.2 | 59.7 3.2 | 42.1 2.2 | 57.1 2.5 |
| 0.5 | 64.1 2.5 | 45.3 2.6 | 62.6 3.1 | |
| 0.8 | 61.1 3.4 | 41.0 1.0 | 59.7 1.1 |
Regarding optimal values the results in Table 8 highlight the influence of the parameter, which balances the contribution of modality-specific and shared representations. These values were selected from the Pareto front obtained by jointly optimizing the two retrieval objectives - mean Recall@10 for 11 and 21 tasks. We observe that moderate values (0.1–0.5) yield the best trade-off between the two objectives, and that the optimal varies across datasets, indicating optimal depends on data distribution.
| Dataset | Mean R@10 (11) | Mean R@10 (21) | |
| MOSI | 0.1 | 21.21 | 22.12 |
| UR-FUNNY | 0.4 | 8.06 | 14.69 |
| MUSTARD | 0.5 | 45.29 | 62.56 |
B.2 Embedding Dimensionality Analysis on the XOR Task
In this experiment, we evaluate how the embedding dimensionality affects each method’s ability to solve the XOR problem for the case of . Figure 6 illustrates the mean accuracy as a function of the embedding dimension. As provided, SYMILE successfully learns to solve the XOR task even with a minimal embedding dimensionality of 8, achieving perfect accuracy thereafter. Our proposed method (ConFu) also manages to solve XOR but requires a larger embedding dimensionality of 64 to reach convergence. In contrast, both TRIANGLE and GRAM fail to solve the task even when the embedding dimensionality is increased up to 1024, implying that the bottleneck lies in their loss formulations rather than in representational capacity.
B.3 Tackling modality shortcuts
Motivated by the hypothesis that the fusion network may exploit shortcut strategies, we investigated the effect of applying dropout at the feature level, specifically, before concatenation and subsequent processing by the fusion MLP. A potential shortcut can arise when the pairwise alignment losses between modalities become sufficiently low, allowing the fusion module to rely predominantly on one modality while suppressing the contribution of the other. In such cases, the network may learn to transmit information through the most predictive modality instead of developing genuinely integrated multimodal representations. Introducing feature-level masking may counteract this behavior by preventing the model from depending exclusively on any single modality and encouraging more balanced multimodal learning.
To examine this effect, we compared model performance with and without masking for multiple values and mask ratios. For each dataset and task, we computed the performance difference across runs and visualized these distributions using boxplots.
As shown in Figure 7, the MOSI dataset consistently benefits from masking across all tasks, classification as well as 11 and 21 retrieval, indicating that feature-level dropout can improve representation robustness and reduce reliance on modality-specific cues. In contrast, the UR-FUNNY dataset exhibits minimal changes, suggesting that masking has limited impact on its multimodal dynamics, while MUStARD experiences performance degradation, likely due to its stronger dependence on precise cross-modal alignment. Moreover, as shown in Table 9, the use of masking in ConFu yields significant performance gains across all querytarget retrieval directions on the MOSI dataset. Overall, these findings suggest that the impact of masking is dataset-dependent: it tends to improve performance when redundant cross-modal information could otherwise enable shortcut learning, but may hinder it when successful alignment relies on using the full representational capacity of each modality.
| Target | Query(s) | Without Mask | With Mask | |
| M1 | M2 | 21.02 1.65 | 25.10 3.51 | +4.08 |
| M3 | 16.41 2.32 | 20.41 1.51 | +4.00 | |
| M23 | 16.73 1.84 | 22.24 3.16 | +5.51 | |
| M2 | M1 | 19.18 1.41 | 25.83 2.98 | +6.65 |
| M3 | 21.02 2.40 | 22.77 0.96 | +1.75 | |
| M13 | 21.63 1.89 | 24.58 2.33 | +2.95 | |
| M3 | M1 | 16.06 1.74 | 23.35 1.61 | +7.29 |
| M2 | 23.47 2.69 | 24.40 1.91 | +0.93 | |
| M12 | 20.50 2.87 | 24.34 3.02 | +3.84 |
B.4 Noise experiments
We investigate the robustness of ConFu under controlled noise-induced distribution shifts applied to either the image or audio modality. Gaussian noise was added at varying severities with standard deviations of 0.05, 0.1, and 0.15 for images, and 0.1, 0.2, and 0.3 for audio. The corresponding SNR values are estimated from the data.
The results in Table 10 show that ConFu exhibits improved robustness to noise compared to competing baselines. Notably, GRAM and TRIANGLE appear largely unaffected when noise is added to audio. However, this behavior is likely attributable to their effective disregard of the audio modality. In contrast, both methods degrade substantially when noise is introduced to the image modality, performing even worse than unimodal TriCLIP.
Overall, ConFu consistently outperforms all baselines across noise levels, including unimodal variants that do not suffer degradation. Its performance trails unimodal TriCLIP-audio slightly only in the extreme case where the visual modality becomes heavily corrupted. A key observation is that Symile, GRAM, and TRIANGLE deteriorate markedly under noise, whereas ConFu incurs only minimal performance loss, particularly when noise is added to audio, where the accuracy drop is negligible. For comparison, TriCLIP experiences roughly a 5% accuracy decline at an audio SNR of 10 dB, while ConFu drops by only 0.2%.
| 10dB SNR | 15dB SNR | 20dB SNR | ||||
| Method | A deg. | V deg. | A deg. | V deg. | A deg. | V deg. |
| Tri-CLIP (V) | 69.0 | 5.5 | 69.0 | 13.2 | 69.0 | 35.4 |
| Tri-CLIP (A) | 26.0 | 31.1 | 29.1 | 31.1 | 30.3 | 31.1 |
| Symile [27] | 58.4 | 21.2 | 59.5 | 27.9 | 60.0 | 40.9 |
| GRAM [6] | 58.9 | 4.0 | 58.3 | 8.94 | 58.1 | 25.9 |
| TRIANGLE [5] | 63.9 | 3.4 | 64.0 | 7.24 | 64.0 | 26.5 |
| ConFu | 71.2 | 30.2 | 71.4 | 33.1 | 71.5 | 45.4 |
B.5 Discussion on Multimodal Competition
In Fig. 8, we present a detailed per-class modality overlap analysis for the SSW60 dataset in the zero-shot classification setting. This experiment visualizes, in the form of a heatmap, the distribution of overlap categories for each label separately. Specifically, each cell indicates the percentage of samples belonging to a given class that fall into one of the following overlap categories: Audiovisual Only, Vision Only, Audio Only, Audiovisual & Vision, Audiovisual & Audio, Vision & Audio, All, or None. These categories capture whether a sample is correctly classified by only one modality, by a specific combination of modalities, by all modalities, or not correctly classified at all.
Higher values in a cell correspond to a larger proportion of samples for which that modality (or modalities) yields a correct prediction. This visualization allows us to inspect class-specific modality behavior, highlighting both complementary and competing contributions across modalities.
We observe that label 31 is exclusively correctly classified by the Audiovisual Only modality, suggesting that neither vision nor audio alone provides sufficient information to recognize this class. This reinforces the need for complementary cross-modal information in order to correctly classify certain bird species.
Appendix C Additional Results
C.1 Few shot adaptation on SSW60, VB100, CUB200 complete results
In this section, we present the complete results of our few-shot adaptation experiments on the SSW60, VB100, and CUB200 [32] datasets. The following figures include all baseline methods for comparison and illustrate adaptation performance across varying number of shot settings. These results provide a comprehensive view of how each method generalizes to limited-data scenarios and adapts to novel classes.
SSW60.
In the multi-frame setting, where mean embeddings over 8 video frames are used, our fusion-based method achieves the best performance. It is important to note that competing approaches do not support multimodal representations, and therefore only unimodal results are available for comparison. In the vision modality, most methods perform similarly, with Symile showing a slight drop. In contrast, in the audio modality, TRIANGLE and GRAM underperform significantly, indicating that their generated audio representations are not informative. In the single-frame setting, ConFu’s audiovisual representations outperform all unimodal variants by a substantial margin.
VB100.
In VB100, audiovisual fusion performs worse than vision-only models. This aligns with our earlier discussion: the audio modality is largely uninformative and even distracting for this dataset, as reflected by the low performance of audio-only baselines.
CUB200.
For CUB200, we evaluate on classes completely unseen during pretraining. Symile achieves the strongest performance, while our method (ConFu) ranks second.
Appendix D Datasets
D.1 Bird-MML Dataset
Dataset Construction.
To construct our dataset, we combined data from three sources: iNaturalist for images, Xeno-Canto for audio, and Wikipedia for textual and class-level grounding. For every bird class present in the VB100 or SSW60 datasets, we collected corresponding images (from iNaturalist), audio recordings (from Xeno-Canto), and Wikipedia articles.
Audio Processing
Audio samples were segmented into 10-second clips (or zero-padded if shorter). We used the tags provided in Xeno-Canto as semantic grounding for the final captions.
Image Captioning
Each image from iNaturalist was processed with InstructBLIP2 to generate visual descriptions, which served as grounding for the final caption.
Model: Salesforce/instructblip-flan-t5-xl
Prompt:
Describe the bird’s colors, size, and shapes.
Wikipedia Text Processing
For each class, we sampled random sections from the corresponding Wikipedia page. These sections were summarized by an LLM to produce textual grounding, generating 100 captions per class. This allowed us to capture diverse factual descriptions of each species.
Model: google/gemma-2-2b-it
Prompt:
You are a naturalist assistant.
Summarize the following text into one coherent caption.
Be concise and factual.
{section_text}
Caption:
Caption Combination & Triplet Construction
In the final step, we matched images, audio clips, and Wikipedia summaries into triplets. To produce a unified caption grounded in all three modalities (image, class, audio), we fed the image caption, audio tags, and a randomly selected Wikipedia caption into an LLM.
Prompt:
You are a naturalist assistant.
Combine the following information into a concise caption (10--20 words).
Include information about the bird’s appearance and sound if available.
Be factual and informative.
Image caption: {image info}
Species caption: {textual info}
Sound tags: {tags}
Caption:
This process results in coherent multimodal captions grounded jointly in visual appearance, species knowledge, and acoustic characteristics.
D.2 Affective Computing Benchmarks
MultiBench [20] is a collection of benchmarking datasets designed to evaluate multimodal representation learning across a wide range of modalities and domains. In our experiments, we used the MOSI, UR-FUNNY, and MUStARD datasets. MultiBench provides pre-extracted features and predefined dataset splits, both of which we adopt in our setup.
MOSI is a multimodal sentiment analysis dataset consisting of 2,199 annotated YouTube video clips. UR-FUNNY contains 16,514 samples extracted from TED Talks and focuses on humor detection in spoken language. MUStARD includes 690 video clips from popular TV shows and is designed for multimodal sarcasm detection.
D.3 AV-MNIST
The AV-MNIST dataset is a multimodal benchmark that pairs degraded visual features with audio spectrograms to evaluate multimodal representation learning. The visual modality consists of 28×28 MNIST digits that have been PCA-projected, retaining only 25% of the total variance. This dimensionality reduction is applied intentionally to weaken the visual signal and encourage effective fusion across modalities.
The audio modality comprises 112×112 spectrograms generated from the Free Spoken Digits Dataset [15], with additive background noise sampled from ESC-50 [25] to increase variability and realism.
The dataset contains 55000 training samples and 10000 test samples. In our experiments, we use the pre-extracted audio spectrograms distributed by MultiBench [20].
Appendix E Generalization to Modalities
For the generalization of our framework to modalities, we extend the objective in Eq. 7 by summing InfoNCE losses over all relevant modality subsets. Specifically, we include all disjoint subset pairs
which correspond to mutual information terms estimated via InfoNCE bounds. This formulation provides a direct extension of our tri-modal framework, allowing contrastive learning to capture both pairwise and higher-order dependencies across arbitrary modality combinations.
To ascertain the above claim, we conducted preliminary experiments in a four-modality setting. This was achieved by splitting the original images into full spatial resolution monochromatic (grayscale) and low-resolution color (RGB), yielding four modalities alongside audio and text without additional data sources. Although artificial, this scenario emulates the behavior of numerous remote sensing platforms (e.g., Pléiades: 0.5 meter monocromatic and 2 meters RGB) and the associated problem of pansharpening. Experiments on VB100, and SSW60 show that the composite objective converges and consistently improves zero-shot performance over a pairwise-only baseline (Table 11), providing evidence that the approach remains effective as additional modalities are incorporated (with VB100 results reflecting the known weakness of the audio signal).
| SSW60 Acc. (%) | VB100 Acc. (%) | |||
| Modality | Pairwise CLIP | ConFu | Pairwise CLIP | ConFu |
| G (grayscale) | 50.33 | 48.46 | 14.48 | 15.47 |
| LR (low-res RGB) | 58.28 | 55.22 | 15.89 | 16.74 |
| A (audio) | 26.67 | 26.02 | 3.04 | 1.77 |
| G + LR | – | 59.39 | – | 18.15 |
| G + A | – | 52.20 | – | 14.05 |
| LR + A | – | 57.63 | – | 15.32 |
| G + LR + A | – | 61.96 | – | 15.32 |
Nevertheless, while the formulation in Eq. 15 captures all possible cross-subset dependencies, its combinatorial nature results in a large number of contrastive terms as increases. In practice, task-specific relaxations of this general objective can substantially reduce computational complexity. For instance 16, when the goal is to retrieve a single target modality, one may restrict the loss to align any subset of the remaining modalities with that target only, i.e., terms of the form . Such relaxations preserve the representational alignment relevant to the retrieval task while avoiding the exponential growth in the number of InfoNCE terms, yielding a more tractable yet principled multimodal contrastive objective.